# Enhancement of non-Gaussianity and nonclassicality of pair coherent states with postselected von Neumann measurement

### v1

@@ III.2 Sum squeezing
where $\Delta V_{\varpi}^{2}=\langle V_{\varpi}^{2}\rangle-\langle V_{\varpi}\rangle^{2}$, and particle number operators for the $a$-mode and $b$-mode are $N_{a}=a^{\dagger}a$ and $N_{b}=b^{\dagger}b$, respectively. If $-1\leq S_{ab}\left(\varpi\right)<0$, the two-mode state exhibits squeezing properties. The more negative the value of $S_{ab}\left(\varpi\right)$, the greater the degree of sum squeezing in the state. The state is maximally squeezed when $S_{ab}\left(\varpi\right)=-1$. By substituting $V_{\varpi}$ into Eq. (25), we obtain the squeezing degree $S_{ab}\left(\varpi\right)$ of normal-ordering operators:
@@
The explicit expression of the SQ degree $S_{ab,\Psi}\left(\varpi\right)$ for the enhanced state $\Psi$ can be derived by substituting the relevant expectation values, as listed in Appendix A. When $\Gamma=0$, no measurements is performed, and the state remains as the initial PCS $|\phi\rangle$. For this trivial case, it is straightforward to verify that the sum squeezing equals to zero, i.e., $S_{ab,\phi}\left(\varpi\right)=0$. This result implies that, regardless of the PND $\delta$ of the two-modes or the angle $\varpi$, the SQ of the initial PCS $|\phi\rangle$ is always zero.
@@
In Fig. 4, we present the SQ degree $S_{ab,\Psi}\left(\varpi\right)$ for the enhanced PCS $|\Psi\rangle$. In Fig. 4(a), we show $S_{ab,\Psi}\left(\varpi\right)$ as a function of the parameter $\gamma$ for different values of the coupling strength parameter $\Gamma$, while fixing the weak value to $\langle\sigma_{x}\rangle_{w}=5.671$, which corresponds to $\alpha=8\pi/9$. As shown in Fig. 4(a), SQ occurs in a small range of the parameter $\gamma$ for appropriate coupling strengths $\Gamma$, provided the weak value of the measured system observable is large. For the anomalous weak value $\langle\sigma_{x}\rangle_{w}=5.671$, that corresponds to $\alpha=8\pi/9$, the minimum value of SQ degree $S_{ab,\Psi}\left(\varpi\right)$ occurs near $\gamma=0.5$ at larger interaction strengths $\Gamma$. This squeezing effect diminishes as $\gamma$ increases beyond this range. To further investigate the dependence of SQ on the weak value, Fig. 4(b) presents the variation $S_{ab,\Psi}\left(\varpi\right)$ as a function of the weak value parameter $\alpha$ for different values of $\Gamma$, while keeping $\gamma=0.5$ fixed. As indicated in Fig. 4(b), the SQ degree $S_{ab,\Psi}\left(\varpi\right)$ takes negative values only for anomalous weak values and non-zero interaction strengths $\Gamma$. Notably, when the weak value angle $\alpha$ approaches $\pi$, the SQ effect becomes more pronounced for suitable $\Gamma$. This numerical result further confirms the amplification effect of weak values in postselected WMs within our scheme.
@@
Figure 4: The SQ of our enhanced PCS $|\Psi\rangle$ after postselected von Neumann measurement. (a) $S_{ab,\Psi}\left(\varpi\right)$ as a function of the parameter $\gamma$ for different coupling strength parameters $\Gamma$, while the weak vale is fixed to $\langle\sigma_{x}\rangle_{w}=5.671$. (b) $S_{ab,\Psi}\left(\varpi\right)$ as a function of the weak value parameter $\alpha$ for different coupling strength parameters $\Gamma$, while the parameter $\gamma$ is fixed at $0.5$. Here, we take $\varpi=0$. Other parameters are the same as in Fig. 2.
@@
In the above discussions, we can conclude that the postselected von Neumann measurement improves various squeezing characteristics of our initial PCS $|\phi\rangle$. That is to say, after postselected von Neumann measurement, the squeezing effects of $|\phi\rangle$ are enhanced for large anomalous weak values of the measured observable, with proper composite system parameters $\gamma$ and $\Gamma$. Besides the squeezing effect, quantum correlation functions and entanglement, which exist between the two modes, also characterize the nature of the PCS $|\phi\rangle$. In the next sections, we investigate the effects of the postselected von Neumann measurement on the quantum correlation functions and entanglement features of the PCS $|\phi\rangle$.
@@ IV.1 Second-order cross-correlation function
As investigated in previous studies (Agarwal, 1988; Hong and Guang-can, 1999), the cross-correlation function $g_{ab,\phi}^{(2)}$ between the $a$-mode and the $b$-mode of the PCS $|\phi\rangle$ decreases as the PND $\delta$ and the parameter $\gamma$ increase, and eventually, $g_{ab,\phi}^{(2)}$ tends to a value of 1, indicating that the correlation between the two modes is almost non-existent.
In Fig. 5 shows the variation of $g_{ab,\Psi}^{(2)}$ for different system parameters. Specifically, in Fig. 5, we plot $g_{a,b,\Psi}^{(2)}$ as a function of the parameter $\gamma$ for different coupling strength parameters $\Gamma$, while fixing the weak value to $\langle\sigma_{x}\rangle_{w}=5.761$. As shown in Fig. 5, the SOCC $g_{ab,\Psi}^{(2)}$ of the state $|\Psi\rangle$ decreases as the coupling strength parameter $\Gamma$ increases. When the parameter $\gamma$ is small, the value of $g_{ab,\Psi}^{(2)}(0)$ is greater than one. However, as $\gamma$ increases, all curves tend to one. The higher the interaction parameter $\Gamma$, the lower the correlation between the two modes of our enhanced PCS $|\Psi\rangle$.
@@
We derived the explicit expressions of $g^{(2)}(0)$ by calculating the average values of the associated operators (see Appendix A). Figure 6 displays the variations of $g_{a,\Psi}^{(2)}(0)$ and $g_{b,\Psi}^{(2)}(0)$ with the state parameter $\gamma$ for different coupling strength parameters $\Gamma$. We fixed the weak value parameter at $\alpha=8\pi/9$ to analyze the effects of large anomalous weak values on the statistical properties of our enhanced state $|\Psi\rangle$. As shown in Fig. 6, the $\Gamma=0$ case corresponds to the $g^{(2)}(0)$ of the initial PCS state $|\phi\rangle$. Its value for both $a$ and $b$-modes is smaller than one and approaches one as the parameter $\gamma$ increases (see green curves in Figs. 6(a) and 6(b)). After the postselected von Neumann measurement ($\Gamma\neq 0$), for small parameter $\gamma$, the values of the second-order correlation functions $g_{a}^{(2)}(0)$ and $g_{b}^{(2)}(0)$ quickly reach their minimum values with increasing coupling strength parameter $\Gamma$, similar to the single-photon state case. On the other hand, for large values of the state parameter ($\gamma\gg 1$), $g_{a}^{(2)}(0)$ and $g_{b}^{(2)}(0)$ become indistinguishable for all coupling strength parameters $\Gamma$ and tend to one, which corresponds to the coherent state case. This result also explains why the SOCC $g_{ab,\Psi}^{(2)}$ of $|\Psi\rangle$ approaches one when $\gamma\gg 1$ (see Fig. 5).
It is interesting to note that when comparing the green thick curve ($\Gamma=0$) with other curves in Figs. 6(a) and 6(b), the second-order correlation function of the $a$-mode and $b$-mode changes more dramatically after the postselected measurement ($\Gamma\neq 0$). Due to the weak value amplification effects, this result allows us to achieve the characteristics of a single-photon state in small regions of the state parameter $\gamma$. Our numerical results suggest that following the postselected von Neumann measurement, the two modes of our enhanced output state $|\Psi\rangle$ could potentially generate a single-photon field (Paul, 1982; Lounis and Orrit, 2005).
Figure 6: Second-order correlation as a function of the $a$ and $b$-modes of the state $|\Psi\rangle$. (a) $g_{a,\Psi}^{(2)}(0)$ as a function of the state parameter $\gamma$ for different coupling strength parameters $\Gamma$ for mode $a$. (b) the $b$ mode’s $g_{b,\Psi}^{(2)}(0)$ as a function of the state parameter $\gamma$ for different coupling strength parameters $\Gamma$ Here, we take weak value as $\langle\sigma_{x}\rangle_{w}=5.761$ and other parameters are the same as in Fig. 2.
@@ V The effects on entanglement
In this section, we investigate the effects of postselected von Neumann measurement on quantum correlations of the PCS. In a previous study (Agarwal and Biswas, 2005), the inseparability of the PCS were confirmed in light of the Peres-Horodecki criterion and various entropies. Here, we separately examine the HZ correlation and EPR correlation to analyze the inseparability of our enhanced output state $|\Psi\rangle$, as defined in Eq. (10).
@@ V.1 HZ correlation
If $E<0$, it implies that the state exhibits entanglement between the two modes. From the CSI, $|\langle ab\rangle|^{2}\leq\langle a^{\dagger}a\rangle\langle b^{\dagger}b\rangle$ (Li et al., 2021), the entanglement condition for two-mode fields is bounded as $-\langle N_{a}\rangle\leq E<0$. It is straightforward to verify that the HZ correlation function $E$ for the initial PCS $|\phi\rangle$ can expressed as:
@@
where $g_{ab,\phi}^{(2)}$ is the SOCC defined in Eq. 28. To evaluate the HZ correlation for our output state $|\Psi\rangle$, we compute $E_{\Psi}$ using Eq. (31), substituting the average values of $\langle ab\rangle,\,\langle a^{\dagger}a\rangle$and $\langle b^{\dagger}b\rangle$ as listed in Appendix A. To demonstrate the effects of postselected von Neumann measurement on the entanglement of the output state $|\Psi\rangle$, we introduce the difference between $E_{\Psi}$ and $E_{\phi}$, defined as $\Delta E=E_{\Psi}-E_{\phi}$. Fig. 7 presents the numerical results for $\Delta E$. If $\Delta E<0$, it indicates that the degree of entanglement has increased compared to the initial state; otherwise, it has decreased. In Fig. 7(a), $\Delta E$ is plotted as a function of the state parameter $\gamma$ for different values of PND denoted by $\delta$, with the coupling strength parameter $\Gamma$ fixed in the WM regime ($\Gamma=0.3$) and a large anomalous weak value $\langle\sigma_{x}\rangle_{w}=5.761$. The results show that $\Delta E$ takes negative values for certain regions of $\gamma$, and these regions expand as $\delta$ increases. This result indicates that the final MD $|\Psi\rangle$ exhibits stronger entangled properties than the initial PCS $|\phi\rangle$ in appropriate parameter regions. To further validate the role of weak value amplification for enhancing the entanglement between two modes of PCS after postselected measurement, Fig. 7(b) illustrates the variation of $\Delta E$ with the weak value parameter $\alpha$ for different values of $\delta$. The curves in Fig. 7(b) reveal that, within the WM regime, if the PND $\delta$ between modes $a$ and $b$ of the PCS is equal to or greater than two (i.e., $\delta\geq 2$), the entanglement between the two modes of the measurement output state $|\Psi\rangle$ surpasses that of the initial state $|\phi\rangle$ for large anomalous weak values.
Figure 7: HZ correlation between two modes of the state $|\Psi\rangle$. (a) $\Delta E$ as a function of the state parameter $\gamma$ for different PND values $\delta$, with $\langle\sigma_{x}\rangle_{w}=5.761$. (b) $\Delta E$ as a function of the weak value parameter $\alpha$ for different PND values $\delta$, with $\gamma=1.5$. Here, we take $\Gamma=0.3$ and other parameters are the same as in Fig. 2.
@@ V.2 EPRs correlation
Here, $Q_{2,\phi}$ is the quadrature squeezing parameter of the initial PCS $|\phi\rangle$ as defined in Eq. (23), and $\Delta F_{2,\phi}^{2}$ represents the variance of the quadrature operator $F_{2}$, defined in Eq. (17), over the state $|\phi\rangle$. For the PCS state $|\phi\rangle$, its EPR correlation is proportional to the variance $\Delta F_{2,\phi}^{2}$. To analyze the effects of the postselected von Neumann measurement on the EPR correlation between the two modes of the state $|\Psi\rangle$, we introduce the difference between $I_{\Psi}$ and $I_{\phi}$, defined as $\Delta I=I_{\Psi}-I_{\phi}$. To observe these effects, we plotted $\Delta I$ as a function of various system parameters, with the numerical results shown in Fig. 8. The variation of $\Delta I$ with changes in the state parameter $\gamma$ for different PND $\delta$ is presented in Fig. 8(a). When the weak value is set to $\langle\sigma_{x}\rangle_{w}=5.761$ and the coupling strength parameter is fixed at $\Gamma=0.3$, $\Delta I$ takes negative values in a range of the state parameter $\gamma$ for cases where $\delta\geq 2$. The regions of negative values become broader and deeper as the PND $\delta$ between the two modes increases. This result implies that in the postselected WM regime ($\Gamma\ll 1$) with large anomalous weak values, the EPR correlation between the two modes of the state $|\Psi\rangle$ is stronger than that of the initial PCS state $|\phi\rangle$ when $\delta\geq 2$. Similar to the HZ correlation case, we examined the dependence of the enhancement of the entanglement on the anomalous weak values. Fig. 8(b) shows $\Delta I$ as a function of the angle $\alpha$. The curves in Fig. 8(b) show that larger PND $\delta$ does not enhance the violation of the EPR correlation for small weak value angles $\alpha$. However, for larger weak value angles $\alpha$, $\Delta I$ becomes negative, and its absolute value increases with increasing PND $\delta$.
@@
Figure 8: EPR correlation differences $\Delta I$ for the states $|\Psi\rangle$ and $|\phi\rangle$. (a) $\Delta I$ as a function of the state parameter $\gamma$ for different PND $\delta$ between the two modes of the state $|\phi\rangle$. (b) $\Delta I$ as a function of the weak value angle $\alpha$ for different PND $\delta$ between the two modes of the state $|\phi\rangle$. Here $\Gamma=0.3$, with (a) $\langle\sigma_{x}\rangle_{w}=5.761$, and (b) $\gamma=1.5$. Other parameters are the same as in Fig. 2.

### v2

@@ III.2 Sum squeezing
where $\Delta V_{\varpi}^{2}=\langle V_{\varpi}^{2}\rangle-\langle V_{\varpi}\rangle^{2}$, and particle number operators for the $a$ mode and $b$ mode are $N_{a}=a^{\dagger}a$ and $N_{b}=b^{\dagger}b$, respectively. If $-1\leq S_{ab}\left(\varpi\right)<0$, the two-mode state exhibits squeezing properties. The more negative the value of $S_{ab}\left(\varpi\right)$, the greater the degree of sum squeezing in the state. The state is maximally squeezed when $S_{ab}\left(\varpi\right)=-1$. By substituting $V_{\varpi}$ into Eq. (25), we obtain the squeezing degree $S_{ab}\left(\varpi\right)$ of normal-ordering operators:
@@
The explicit expression of the SQ degree $S_{ab,\Psi}\left(\varpi\right)$ for the final MD state $|\Psi\rangle$ can be derived by substituting the relevant expectation values, as listed in Appendix A. When $\Gamma=0$, no measurements is performed, and the state remains as the initial PCS $|\phi\rangle$. For this trivial case, it is straightforward to verify that the sum squeezing equals to zero, i.e., $S_{ab,\phi}\left(\varpi\right)=0$. This result implies that, regardless of the PND $\delta$ of the two-modes or the angle $\varpi$, the SQ of the initial PCS $|\phi\rangle$ is always zero.
@@
In Fig. 4, we present the SQ degree $S_{ab,\Psi}\left(\varpi\right)$ for the final MD state $|\Psi\rangle$. In Fig. 4(a), we show $S_{ab,\Psi}\left(\varpi\right)$ as a function of the state parameter $\gamma$ for different values of the coupling strength parameter $\Gamma$, while fixing the weak value to $\langle\sigma_{x}\rangle_{w}=5.671$, which corresponds to $\alpha=8\pi/9$. As shown in Fig. 4(a), SQ occurs in a small range of the state parameter $\gamma$ for appropriate coupling strengths $\Gamma$, provided the weak value of the measured system observable is large. For the anomalous weak value $\langle\sigma_{x}\rangle_{w}=5.671$, that corresponds to $\alpha=8\pi/9$, the minimum value of SQ degree $S_{ab,\Psi}\left(\varpi\right)$ occurs near $\gamma=0.5$ at larger coupling strength parameters $\Gamma$. This squeezing effect diminishes as $\gamma$ increases beyond this range. To further investigate the dependence of SQ on the weak value, Fig. 4(b) presents the variation $S_{ab,\Psi}\left(\varpi\right)$ as a function of the weak value parameter $\alpha$ for different values of $\Gamma$, while keeping $\gamma=0.5$ fixed. As indicated in Fig. 4(b), the SQ degree $S_{ab,\Psi}\left(\varpi\right)$ takes negative values only for anomalous weak values and non-zero coupling strength parameters $\Gamma$. Notably, when the weak value angle $\alpha$ approaches $\pi$, the SQ effect becomes more pronounced for suitable coupling strength parameters $\Gamma$. This numerical result further confirms the amplification effect of weak values in postselected WMs within our scheme.
@@
Figure 4: The SQ of the final MD state $|\Psi\rangle$ under postselected von Neumann measurement. (a) $S_{ab,\Psi}\left(\varpi\right)$ as a function of the state parameter $\gamma$ for different coupling strength parameters $\Gamma$, while the weak vale is fixed to $\langle\sigma_{x}\rangle_{w}=5.671$. (b) $S_{ab,\Psi}\left(\varpi\right)$ as a function of the weak value parameter $\alpha$ for different coupling strength parameters $\Gamma$, while the state parameter $\gamma$ is fixed at $0.5$. Here, we take $\varpi=0$. Other parameters are the same as in Fig. 2.
@@
In the above discussions, we can conclude that the postselected von Neumann measurement improves various squeezing characteristics of our initial PCS $|\phi\rangle$. That is to say, after postselected von Neumann measurement, the squeezing effects of $|\phi\rangle$ are enhanced for large anomalous weak values of the measured observable, with proper composite state parameters $\gamma$ and coupling strength parameters $\Gamma$. Besides the squeezing effect, quantum correlation functions and entanglement, which exist between the two modes, also characterize the nature of the PCS $|\phi\rangle$. In the next sections, we investigate the effects of the postselected von Neumann measurement on the quantum correlation functions and entanglement features of the PCS $|\phi\rangle$.
@@ IV.1 Second-order cross-correlation function
As investigated in previous studies (Agarwal, 1988; Hong and Guang-can, 1999), the cross-correlation function $g_{ab,\phi}^{(2)}$ between the $a$-mode and the $b$-mode of the PCS $|\phi\rangle$ decreases as the PND $\delta$ and the state parameter $\gamma$ increase, and eventually, $g_{ab,\phi}^{(2)}$ tends to a value of 1, indicating that the correlation between the two modes is almost non-existent.
In Fig. 5 shows the variation of $g_{ab,\Psi}^{(2)}$ for different system parameters. Specifically, in Fig. 5, we plot $g_{a,b,\Psi}^{(2)}$ as a function of the parameter $\gamma$ for different coupling strength parameters $\Gamma$, while fixing the weak value to $\langle\sigma_{x}\rangle_{w}=5.761$. As shown in Fig. 5, the SOCC $g_{ab,\Psi}^{(2)}$ of the state $|\Psi\rangle$ decreases as the coupling strength parameter $\Gamma$ increases. When the state parameter $\gamma$ is small, the value of $g_{ab,\Psi}^{(2)}(0)$ is greater than one. However, as $\gamma$ increases, all curves tend to one. The higher the coupling strength parameters $\Gamma$, the lower the correlation between the two modes of the final MD state $|\Psi\rangle$.
@@
We derived the explicit expressions of $g^{(2)}(0)$ by calculating the average values of the associated operators (see Appendix A). Figure 6 displays the variations of $g_{a,\Psi}^{(2)}(0)$ and $g_{b,\Psi}^{(2)}(0)$ with the state parameter $\gamma$ for different coupling strength parameters $\Gamma$. We fixed the weak value parameter at $\alpha=8\pi/9$ to analyze the effects of large anomalous weak values on the statistical properties of the final MD state $|\Psi\rangle$. As shown in Fig. 6, the $\Gamma=0$ case corresponds to the $g^{(2)}(0)$ of the initial PCS state $|\phi\rangle$. Its value for both $a$ and $b$-modes is smaller than one and approaches one as the state parameter $\gamma$ increases (see green curves in Figs. 6(a) and 6(b)). After the postselected von Neumann measurement ($\Gamma\neq 0$), for small parameter $\gamma$, the values of the second-order correlation functions $g_{a}^{(2)}(0)$ and $g_{b}^{(2)}(0)$ quickly reach their minimum values with increasing coupling strength parameter $\Gamma$, similar to the single-photon state case. On the other hand, for large values of the state parameter ($\gamma\gg 1$), $g_{a}^{(2)}(0)$ and $g_{b}^{(2)}(0)$ become indistinguishable for all coupling strength parameters $\Gamma$ and tend to one, which corresponds to the coherent state case. This result also explains why the SOCC $g_{ab,\Psi}^{(2)}$ of $|\Psi\rangle$ approaches one when $\gamma\gg 1$ (see Fig. 5).
It is interesting to note that when comparing the green thick curve ($\Gamma=0$) with other curves in Figs. 6(a) and 6(b), the second-order correlation function of the $a$-mode and $b$-mode changes more dramatically after the postselected measurement ($\Gamma\neq 0$). Due to the weak value amplification effects, this result allows us to achieve the characteristics of a single-photon state in small regions of the state parameter $\gamma$. Our numerical results suggest that following the postselected von Neumann measurement, the two modes of the final MD state $|\Psi\rangle$ could potentially generate a single-photon field (Paul, 1982; Lounis and Orrit, 2005).
Figure 6: Second-order correlation as a function of the $a$ and $b$ modes of the final MD state $|\Psi\rangle$. (a) $g_{a,\Psi}^{(2)}(0)$ as a function of the state parameter $\gamma$ for different coupling strength parameters $\Gamma$ for mode $a$. (b) the $b$ mode’s $g_{b,\Psi}^{(2)}(0)$ as a function of the state parameter $\gamma$ for different coupling strength parameters $\Gamma$ Here, we take weak value as $\langle\sigma_{x}\rangle_{w}=5.761$ and other parameters are the same as in Fig. 2.
@@ V The effects on entanglement
In this section, we investigate the effects of postselected von Neumann measurement on quantum correlations of the PCS. In a previous study (Agarwal and Biswas, 2005), the inseparability of the PCS were confirmed in light of the Peres-Horodecki criterion and various entropies. Here, we separately examine the HZ correlation and EPR correlation to analyze the inseparability of the final MD state $|\Psi\rangle$, as defined in Eq. (10).
@@ V.1 HZ correlation
If $E<0$, it implies that the state exhibits entanglement between the two modes. From the CSI, $|\langle ab\rangle|^{2}\leq\langle a^{\dagger}a\rangle\langle bb^{\dagger}\rangle$ (Li et al., 2021), the entanglement condition for two-mode fields characterized by HZ correlation is bounded as $-\langle N_{a}\rangle\leq E<0$. It is straightforward to verify that the HZ correlation function $E$ for the initial PCS $|\phi\rangle$ can expressed as:
@@
where $g_{ab,\phi}^{(2)}$ is the SOCC defined in Eq. 28. To evaluate the HZ correlation for our output state $|\Psi\rangle$, we compute $E_{\Psi}$ using Eq. (31), substituting the average values of $\langle ab\rangle,\,\langle a^{\dagger}a\rangle$and $\langle b^{\dagger}b\rangle$ as listed in Appendix A. To demonstrate the effects of postselected von Neumann measurement on the entanglement of the output state $|\Psi\rangle$, we introduce the difference between $E_{\Psi}$ and $E_{\phi}$, defined as $\Delta E=E_{\Psi}-E_{\phi}$. Fig. 7 presents the numerical results for $\Delta E$. If $\Delta E<0$, it indicates that the degree of entanglement has increased compared to the initial state; otherwise, it has decreased. In Fig. 7(a), $\Delta E$ is plotted as a function of the state parameter $\gamma$ for different values of PND denoted by $\delta$, with the coupling strength parameter $\Gamma$ fixed in the WM regime ($\Gamma=0.3$) and a large anomalous weak value $\langle\sigma_{x}\rangle_{w}=5.761$. The results show that $\Delta E$ takes negative values for certain regions of $\gamma$, and these regions expand as $\delta$ increases. This result indicates that the final MD state $|\Psi\rangle$ exhibits stronger entangled properties than the initial PCS $|\phi\rangle$ in appropriate parameter regions. To further validate the role of weak value amplification for enhancing the entanglement between two modes of PCS after postselected measurement, Fig. 7(b) illustrates the variation of $\Delta E$ with the weak value parameter $\alpha$ for different values of $\delta$. The curves in Fig. 7(b) reveal that, within the WM regime, if the PND $\delta$ between modes $a$ and $b$ of the PCS is equal to or greater than two (i.e., $\delta\geq 2$), the entanglement between the two modes of the measurement output state $|\Psi\rangle$ surpasses that of the initial state $|\phi\rangle$ for large anomalous weak values.
Figure 7: HZ correlation between two modes of the final MD state $|\Psi\rangle$. (a) $\Delta E$ as a function of the state parameter $\gamma$ for different PND values $\delta$, with $\langle\sigma_{x}\rangle_{w}=5.761$. (b) $\Delta E$ as a function of the weak value parameter $\alpha$ for different PND values $\delta$, with $\gamma=1.5$. Here, we take $\Gamma=0.3$ and other parameters are the same as in Fig. 2.
@@ V.2 EPRs correlation
Here, $Q_{2,\phi}$ is the quadrature squeezing parameter of the initial PCS $|\phi\rangle$ as defined in Eq. (23), and $\Delta F_{2,\phi}^{2}$ represents the variance of the quadrature operator $F_{2}$, defined in Eq. (17), over the state $|\phi\rangle$. For the PCS state $|\phi\rangle$, its EPR correlation is proportional to the variance $\Delta F_{2,\phi}^{2}$. To analyze the effects of the postselected von Neumann measurement on the EPR correlation between the two modes of the state $|\Psi\rangle$, we introduce the difference between $I_{\Psi}$ and $I_{\phi}$, defined as $\Delta I=I_{\Psi}-I_{\phi}$. To observe these effects, we plotted $\Delta I$ as a function of various system parameters, with the numerical results shown in Fig. 8. The variation of $\Delta I$ with changes in the state parameter $\gamma$ for different PND $\delta$ is presented in Fig. 8(a). When the weak value is set to $\langle\sigma_{x}\rangle_{w}=5.761$ and the coupling strength parameter is fixed at $\Gamma=0.3$, $\Delta I$ takes negative values in a range of the state parameter $\gamma$ for cases where $\delta\geq 2$. The regions of negative values become broader and deeper as the PND $\delta$ between the two modes increases. This result implies that in the postselected WM regime ($\Gamma<1$) with large anomalous weak values, the EPR correlation between the two modes of the final MD state $|\Psi\rangle$ is stronger than that of the initial PCS state $|\phi\rangle$ when $\delta\geq 2$. Similar to the HZ correlation case, we examined the dependence of the enhancement of the entanglement on the anomalous weak values. Fig. 8(b) shows $\Delta I$ as a function of the angle $\alpha$. The curves in Fig. 8(b) show that larger PND $\delta$ does not enhance the violation of the EPR correlation for small weak value angles $\alpha$. However, for larger weak value angles $\alpha$, $\Delta I$ becomes negative, and its absolute value increases with increasing PND $\delta$.
@@
Figure 8: EPR correlation differences $\Delta I$ for the final MD state $|\Psi\rangle$ and initial PCS $|\phi\rangle$. (a) $\Delta I$ as a function of the state parameter $\gamma$ for different PND $\delta$ between the two modes of the state $|\phi\rangle$. (b) $\Delta I$ as a function of the weak value angle $\alpha$ for different PND $\delta$ between the two modes of the state $|\phi\rangle$. Here $\Gamma=0.3$, with (a) $\langle\sigma_{x}\rangle_{w}=5.761$, and (b) $\gamma=1.5$. Other parameters are the same as in Fig. 2.